OpenStudy (anonymous):

Find the point on the terminal side of θ =-3pie/4 that has an x coordinate of -1

OpenStudy (anonymous):

because x=-1 that's why tan is -1?

OpenStudy (anonymous):

Can somebody please explain >.<

OpenStudy (anonymous):

is it because x coordinate is -1? so it would be the opposite of 3pie/4?

OpenStudy (anonymous):

let me rephrase that, I understand the x coordinate of -1, I know that 3pie/4 is in quadrant 3, what throws me off is the negative sign in front of it.

OpenStudy (anonymous):

okay...?

OpenStudy (anonymous):

\[\text{Tan}\left[\frac{-3\pi }{4}\right]\text{=}\text{Tan}[-135 {}^{\circ}]=1 \]The point of interest is (-1,-1)

OpenStudy (anonymous):

okay, thank you so much :)

OpenStudy (anonymous):

You're very welcome.